# R-Bot: An LLM-based Query Rewrite System

### v1

@@ Abstract.
Query rewrite is essential for optimizing SQL queries to improve their execution efficiency without changing their results. Traditionally, this task has been tackled through heuristic and learning-based methods, each with its limitations in terms of inferior quality and low robustness. Recent advancements in LLMs offer a new paradigm by leveraging their superior natural language and code comprehension abilities. Despite their potential, directly applying LLMs like GPT-4 has faced challenges due to problems such as hallucinations, where the model might generate inaccurate or irrelevant results. To address this, we propose R-Bot, an LLM-based query rewrite system with a systematic approach. We first design a multi-source rewrite evidence preparation pipeline to generate query rewrite evidences for guiding LLMs to avoid hallucinations. We then propose a hybrid structure-semantics retrieval method that combines structural and semantic analysis to retrieve the most relevant rewrite evidences for effectively answering an online query. We next propose a step-by-step LLM rewrite method that iteratively leverages the retrieved evidences to select and arrange rewrite rules with self-reflection. We conduct comprehensive experiments on widely used benchmarks, and demonstrate the superior performance of our system, R-Bot, surpassing state-of-the-art query rewrite methods.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Query rewrite is designed to transform an SQL query into a logically equivalent version that is more efficient to execute, playing a crucial role in enhancing query performance in numerous practical scenarios. Despite its significance, the process of query rewrite is NP-hard (Zhou et al., 2021, 2023), meaning there is a vast collection of possible rewrite rules, and the number of potential rule combinations increases exponentially. This complexity makes identifying an effective combination of rules a challenging and laborious task. There are two main paradigms for addressing this challenge.
Heuristic-Based Methods. Some heuristic-based methods apply the rules in a fixed order derived from practical experience (e.g., PostgreSQL (pos, 2024)). However, they may not achieve optimal results for queries requiring different rule orders, thereby risking the omission of essential rewrite sequences. Furthermore, other heuristic-based methods (e.g., Volcano (Graefe and McKenna, 1993)) attempt to comprehensively explore various rule orders through heuristic acceleration. Nevertheless, they might overlook dependencies among rules, such as an initially inapplicable rule could be activated by another, leading to the potential neglect of vital rewrite sequences. Consequently, heuristic-based approaches are often criticized for their inferior quality.
Learning-Based Methods. To optimize query rewrite, learning-based methods have been proposed (Zhou et al., 2021, 2023). These methods employ neural networks that are trained on historical query rewrites to identify and apply the most advantageous rules for rewriting a query. However, learning-based approaches face criticism for their low robustness. For example, models trained through these methods struggle to adapt to unseen database schemas without undergoing additional training on hundreds of new query rewrite examples, which may not be readily available in many scenarios.
Recent advances in large language models (LLMs) have shown superiority in understanding natural language and code, as well as reasoning ability (Hu et al., 2024; Achiam et al., 2023; Du et al., 2024; Imani et al., 2023). As LLMs can capture the query rewrite capabilities by pre-training from database forums and codes, encompassing both the direct rewriting ability of applying a rule and the indirect rewriting ability to draw inspiration from multiple dependent rules, we can leverage LLMs to guide the query rewrite.
Table 1. R-Bot vs Existing Query Rewrite Methods.
@@
To realize this ambition, we aim to develop an LLM-based query rewrite system with two main advantages: $(i)$ High Quality. On one hand, our system can figure out the potential rules indirectly related to the query (e.g., ones that become applicable only after applying other rules). On the other hand, our system can understand the interrelations among rules and generate an effective sequence of rules for holistic improvements. $(ii)$ High Robustness. Compared to conventional learning-based approaches, our system using LLMs benefits from extensive pre-training, endowing it with superior adaptability and robustness.
@@
C2: How to mitigate LLM’s faithfulness hallucination in query rewrite? LLM also encounters challenges in accurately analyzing complex queries, such as those containing multiple sub-queries, and in fully leveraging detailed rewrite evidences. Consequently, it becomes essential to develop a multi-step LLM rewrite algorithm that breaks down the query rewrite process into more manageable segments, thereby aligning with LLM’s capabilities for better performance.
@@
To tackle the above challenges, we propose R-Bot, an LLM-based query rewrite system designed with a systematic approach. First, we gather and prepare rewrite evidences from diverse sources, including integrated rewrite rules aggregating from rewrite documents and summarizing from the complex rule codes, as well as high-quality Q&As from database forums (addressing C1). Second, for an input SQL query, we propose a hybrid structure-semantics method for retrieving pertinent evidences, including rewrite rules by matching functions and rewrite Q&As with both query structure and rewrite semantics similarities (addressing C1). To enable LLM to comprehend the rewrite evidences, we synthesize this information into rewrite recipes, which detail in natural language how to utilize the Q&As and rewrite rules for rewriting an SQL query, and we retrieve most relevant rewrite recipes to guide LLMs for query rewrite (addressing C2). Third, we design a step-by-step LLM rewrite algorithm, which guides LLM to iteratively utilize the rewrite recipes to refine its rewrite rule selection and ordering, possibly choosing a more promising rule sequence by reflecting and self-improving the rewrite process (addressing C2).
@@
$(2)$ We design a multi-source rewrite evidence preparation pipeline, including clustering-based document re-organization and hierarchical code summarization for rewrite codes (see Section 4).
@@
$(5)$ Our experimental results on widely used benchmarks demonstrate that R-Bot can significantly outperform existing state-of-the-art query rewrite methods (see Section 7).
@@ 2. Preliminaries
In this section, we first formalize the problem of query rewrite (see Section 2.1). Then we introduce fundamental LLM concepts used in R-Bot (see Section 2.2).
@@ 2.1. Query Rewrite
Consider a query $q$ and a set of rewrite rules $R$. Assume $\alpha$ is a sequence of certain rewrite rules selected from $R$. Let $q^{\alpha}$ denote the rewritten query by sequentially applying the rules in $\alpha$ to rewrite the query $q$. The objective of query rewrite is to obtain an optimal rule sequence $\alpha^{*}$, such that the execution cost of $q^{\alpha^{*}}$ is minimized among all possible rewritten queries of $q$.
The query rewrite problem has been proven to be NP-hard (Zhou et al., 2021). Traditional methods cannot select high-quality rules. To address this limitation, we advocate for utilizing LLM to select rewrite rules. To address the hallucination problem of LLMs (Huang et al., 2023), we perform an offline stage to extract query rewrite evidences, including rewrite Q&As and rewrite rule specifications, and store them as Q&A repository and rewrite rule specification repository respectively. During the online phase, given an input SQL query, we retrieve relevant Q&As and rule specifications, and generate rewrite recipes, which outline how to rewrite a query using rewrite Q&As and rule specifications, thereby assisting LLMs in comprehending the rewrite evidences. With the assistance of the rewrite recipe, LLMs are then guided through a step-by-step process to judiciously select and apply rewrite rules to the query. Next we formally define these notations.
@@
For instance, Figure 1 shows examples of rule specification recipe and Q&A recipe.
@@ 2.2. Large Language Models
We introduce some fundamental concepts of LLMs, including LLM prompting and retrieval augmented generation (RAG), which are pivotal in composing an evidence-driven query rewrite system.
LLM Prompting. After instruction tuning on large corpus of human interaction scenarios, LLM is capable of following human instructions to complete different tasks (Ouyang et al., 2022; Dong et al., 2022). Specifically, given the task input $X$, we concatenate a prompt $P$, denoted as $X^{\prime}=P\bigoplus X$, which represents a comprehensive task description. LLM then fulfills the task, and generates the task output. Note that pre-training on trillions of bytes of data enables LLM to generalize to diverse tasks by simply adjusting the prompt  (Achiam et al., 2023; Khattab et al., 2023).
Retrieval Augmented Generation. During completing tasks with prompting, LLMs often generate erroneous or meaningless responses, i.e., the hallucination problem (Huang et al., 2023). To mitigate the problem, retrieval augmented generation (RAG) is utilized as an advanced LLM prompting technique, which is composed of three steps (Gao et al., 2023): $(i)$ indexing, which prepares chunks of the task-specific knowledge source with embedding vector index; $(ii)$ retrieval, given a user query, using the index to retrieve the most relevant chunks to the query; $(iii)$ generation, guiding LLM to generate answers with the retrieved chunks as a prompt.
@@
However, existing RAG and prompting techniques face limitations when applied directly to query rewrite for two primary reasons. First, there is absence of embedding methods capable of accurately evaluating the similarity between Q&As and the input query. For example, the conventional RAG technique relies on text embeddings for both the Q&A and the query, capturing only their semantic similarity while neglecting crucial structural information pertinent to query rewrite. To address this gap, we propose a hybrid structure-semantics approach for retrieving relevant Q&As, as detailed in Section 5. Second, the multitude of rules presents a significant challenge for LLMs, as directly arranging these rules can lead to serious hallucination problems. This necessitates the development of a task-specific, step-by-step LLM algorithm that breaks down the query rewrite process into simpler, more manageable stages, as elaborated in Section 6.
@@ 3. The Overview of R-Bot
Our R-Bot system includes an offline stage and an online stage, as described in Figure 2.
Offline Rewrite Evidence Preparation. This stage aims to extract rewrite Q&As from the Web and rewrite rule specifications from rewrite codes and database documents, subsequently storing them in Q&A repository and rule specification repository, respectively. These resources will be utilized to assist in rewriting an online SQL query. Note that rewrite rule specifications are more general that can be applied to multiple SQL queries. In contrast, rewrite Q&As are usually specific and are applied to a particular SQL query.
@@
Additionally, we construct efficient indexes to enhance the performance of Q&A and rule specification retrieval. We will discuss the detailed techniques in Section 4.
Online LLM-Guided Query Rewrite. Given an online SQL query, this stage retrieves pertinent Q&As and rule specifications, leverages them to rewrite the query, and offers reflections that not only complete the rewrite process but facilitate its further refinement.
(1) Hybrid Structure-Semantic Evidence Retrieval. For an online SQL query, we retrieve relevant evidences from rewrite rule specification repository and Q&A repository.
  1. (i)
    Rule specification retrieval. As the rule specification has matching conditions and matching functions, we can easily retrieve the relevant rule specifications whose matching functions are satisfied by the input SQL query;
  2. (ii)
    Q&A retrieval. There are two types of Q&As that could potentially be relevant to the SQL query. First, the SQL questioned in Q&A structurally matches the input SQL query. We require a structure-aware matching method to retrieve such Q&As. To this end, we propose a query structure embedding composed of $(i)$ the query template embedded by pre-trained embedding, and $(ii)$ one-hot embedding which represents the query’s matched rule specifications to assess the structural similarities. Second, Q&A semantically matches the input SQL query, such as the rewrite explanations provided in the answer section of the Q&A. However, SQL usually has no natural language rewrite explanations. To this end, we can leverage the retrieved rule specification to identify relevant Q&As based on their semantics. We propose a semantics embedding method to embed the rule specification and Q&A, and then assess their similarities based on a similarity function (e.g., $L^{2}$-distance (chr, 2024)). Lastly, to retrieve Q&As relevant to the input query both structurally and semantically, we merge the query’s structural embedding with its semantic embedding into a unified representation. We then build an embedding index for Q&A repository offline. For an online query, we utilize this index to efficiently retrieve top-$k$ relevant Q&As.
@@
  1. (i)
    Q&A-based rule selection and ordering. Given the rewrite rules and a sorted sequence of Q&As retrieved from the previous step, ranked by the relevance score, we select and rank the rules. We first initialize the score of each rule as 0. Then for each pair of a rule and a Q&A, we use LLM to evaluate whether they are relevant, i.e., whether the rule is applicable in light of the Q&A. If applicable, we increase the score of the rule by the score of this Q&A. Then by enumerating all the pairs of rules and Q&As, we can get the final score of each rule and rank the rules based on the final score. Since the relevance can be evaluated by LLM offline, this step can be efficiently executed by algorithms.
  2. (ii)
    Recipe-based rule selection and ordering. Building on the preliminary rule sequence established in the previous step, we utilize LLM to sift through and exclude any rules that do not align with the recipe. A straightforward way is to enumerate each pair of recipes returned by step (1) and rules returned by step $(2.i)$, and ask LLM to evaluate their relevance and rank the rules. However, there are two limitations. First, it may overlook the rule relevance, since a recipe may encompass multiple rules. Second, since the recipe is generated from queries and cannot be evaluated offline, assessing each pair with LLM becomes costly. To address this issue, we propose an effective filtering method to efficiently select the rules.
  3. (iii)
    Rule-based rewrite. Based on the selected rules, we input them to the query rewrite engine to rewrite the query.
@@
(3) Rewrite Reflection. It provides rewrite reflections to either further refine the query or to finalize the rewrite process. It has two reflection resources. The first involves getting the cost of the rewritten query from databases, comparing it with the cost of the query prior to its rewrite, and returning complete if the cost of the rewritten query is smaller; continue otherwise. The second involves asking LLMs to check whether or not all the rewrite recipes are realized by the query rewrite in this step, and returning complete if yes; continue otherwise. So there are four possible reflections.
  1. (i)
    complete, complete: It finalizes the rewrite process and returns the rewritten query.
  2. (ii)
    complete, continue: It further rewrites the query by jumping to step (1) with the previously rewritten query as the input. Specifically, it starts a new round of LLM-guided query rewrite, where new rewrite evidences can be retrieved and new rules can be selected based on the previously rewritten query.
  3. (iii)
    continue, complete: It further refines the query by jumping to step $(2.ii)$, focusing on reordering the existing set of rules. Specifically, besides the recipes and the rules, we further input the rules actually used in the previous rewrite process, and instruct LLM to prioritize the unused rules.
  4. (iv)
    continue, continue: This approach integrates elements from branches $(3.ii)$ and $(3.iii)$. Initially, it defaults to branch $(3.iii)$ unless this path is revisited excessively, surpassing a predefined threshold. Under these circumstances, given that branch $(3.iii)$ has exhaustively explored the rules for the current query, the process transitions to branch $(3.ii)$. This shift initiates a new cycle, aiming to further refine the query.

### v2

@@ Abstract.
Query rewrite is essential for optimizing SQL queries to improve their execution efficiency without changing their results. Traditionally, this task has been tackled through heuristic and learning-based methods, each with its limitations in terms of inferior quality and low robustness. Recent advancements in LLMs offer a new paradigm by leveraging their superior natural language and code comprehension abilities. Despite their potential, directly applying LLMs like GPT-4 has faced challenges due to problems such as hallucinations, where the model might generate inaccurate or irrelevant results. To address this, we propose R-Bot, an LLM-based query rewrite system with a systematic approach. We first design a multi-source rewrite evidence preparation pipeline to generate query rewrite evidences for guiding LLMs to avoid hallucinations. We then propose a hybrid structure-semantics retrieval method that combines structural and semantic analysis to retrieve the most relevant rewrite evidences for effectively answering an online query. We next propose a step-by-step LLM rewrite method that iteratively leverages the retrieved evidences to select and arrange rewrite rules with self-reflection. We conduct comprehensive experiments on real-world datasets and widely used benchmarks, and demonstrate the superior performance of our system, R-Bot, surpassing state-of-the-art query rewrite methods. The R-Bot system has been deployed at Huawei and with real customers, and the results show that the proposed R-Bot system achieves lower query latency.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/curtis-sun/LLM4Rewrite.
@@ 1. Introduction
Query rewrite is designed to transform an SQL query into a logically equivalent version that is more efficient to execute, playing a crucial role in enhancing query performance in numerous practical scenarios. In Huawei’s real-world database deployments, slow queries are regularly identified and optimized for improved performance. For instance, while migrating an enterprise’s core application, we rewrote 20 critical queries, leading to a 3.7x reduction in workload latency. Despite its significance, the process of query rewrite is NP-hard (Zhou et al., 2021, 2023), meaning there is a vast collection of possible rewrite rules, and the number of potential rule combinations increases exponentially. This complexity makes identifying an effective combination of rules a challenging and laborious task. There are two main paradigms for addressing this challenge.
Table 1. R-Bot v.s. Existing Query Rewrite Methods.
@@
Heuristic-based Methods. Some heuristic-based methods apply the rules in a fixed order derived from practical experience (e.g., PostgreSQL (pos, 2024)). However, they may not achieve optimal results for queries requiring different rule orders, thereby risking the omission of essential rewrite sequences. Furthermore, other heuristic-based methods (e.g., Volcano (Graefe and McKenna, 1993)) attempt to comprehensively explore various rule orders through heuristic acceleration. Nevertheless, they might overlook dependencies among rules, such as an initially inapplicable rule could be activated by another, leading to the potential neglect of vital rewrite sequences. Thus, heuristic-based approaches are often criticized for their inferior quality.
Learning-based Methods. To optimize query rewrite, learning-based methods have been proposed (Zhou et al., 2021, 2023). These methods employ neural networks that are trained on historical query rewrites to identify and apply the most advantageous rules for rewriting a query. However, learning-based approaches face criticism for their low robustness. For example, models trained through these methods struggle to adapt to unseen database schemas without undergoing additional training on new query rewrite examples (e.g., hundreds of examples), which may not be readily available in real scenarios.
Recent advances in large language models (LLMs) have shown superiority in understanding natural language and code, as well as reasoning ability (Hu et al., 2024; Achiam et al., 2023; Du et al., 2024; Imani et al., 2023; Zheng et al., 2024; Liu et al., 2024b; Fan et al., 2024; Li et al., 2024a; Zhou et al., 2025b; Li et al., 2025; Wang et al., 2025; Wang and Li, 2025; Zhou et al., 2025a; Sun et al., 2025a). As LLMs can capture the query rewrite capabilities by pre-training from database forums and codes, encompassing both the direct rewriting ability of applying a rule and the indirect rewriting ability to draw inspiration from multiple dependent rules, we can leverage LLMs to guide the query rewrite, particularly for slow queries that often remain as bottlenecks.
To realize this target, we aim to develop an LLM-based query rewrite system with three main advantages: (1) High Quality. On one hand, our system can figure out the potential rules that have implicit relations with the query (e.g., ones that become applicable only after applying other rules). On the other hand, our system can understand the interrelations among rules and generate an effective sequence of rules for holistic improvements. (2) Zero-Shot Robustness. Unlike conventional learning-based approaches that are limited to in-distribution data (Zhou et al., 2021, 2023), our system harnesses LLMs, whose extensive pre-training empowers it to adapt to new datasets seamlessly without the need for additional retraining. (3) Executability and Equivalence. Our system ensures the rewritten query is both executable and functionally equivalent to the original one, as it performs query rewrite by selecting and ordering well-crafted rewrite rules from established query optimization engines, rather than directly rewriting the queries via LLMs.
@@
C2: How to mitigate LLM’s faithfulness hallucination in query rewrite? LLM encounters challenges in accurately analyzing complex queries, such as those containing multiple sub-queries, and in fully leveraging detailed rewrite evidences. Thus, it becomes essential to develop a multi-step LLM rewrite method that breaks down the query rewrite process into more manageable segments, thereby aligning with LLM’s capabilities for better performance.
@@
To tackle the above challenges, we propose R-Bot, an LLM-based query rewrite system designed with a systematic approach. First, we gather and prepare rewrite evidences from diverse sources, including integrated well-formatted rewrite rules aggregating from rewrite documents and summarizing from the complex rule codes, as well as high-quality Q&As from database forums (addressing C1). Second, for an input SQL query, we propose a hybrid structure-semantics method for retrieving pertinent evidences, including rewrite rules by matching functions and rewrite Q&As with both query structure and rewrite semantics similarities (addressing C1). To enable LLM to comprehend the rewrite evidences, we synthesize this information into rewrite recipes, which detail in natural language how to utilize the Q&As and rewrite rules for rewriting an SQL query, and we retrieve most relevant rewrite recipes to guide LLMs for query rewrite (addressing C2). Third, we design a step-by-step LLM rewrite algorithm, which guides LLM to iteratively utilize the rewrite recipes to refine its rewrite rule selection and ordering, possibly choosing a more promising rule sequence by reflecting and self-improving the rewrite process (addressing C2).
@@
$(2)$ We design a multi-source rewrite evidence preparation pipeline, including clustering-based document reorganization and hierarchical code summarization for rewrite codes (see Section 4).
@@
$(5)$ Our experimental results on real-world datasets and widely used benchmarks demonstrate that R-Bot can significantly outperform existing state-of-the-art query rewrite methods. We have deployed R-Bot at Huawei with real customers, achieving much higher real-world query rewrite performance (see Section 7).
@@ 2.1. Query Rewrite
Consider a query $q$ and a set of rewrite rules $R$. Assume $\alpha$ is a sequence of certain rewrite rules selected from $R$. Let $q^{\alpha}$ denote the rewritten query by sequentially applying the rules in $\alpha$ to rewrite $q$. Query rewrite aims to obtain an optimal rule sequence $\alpha^{*}$, such that the execution cost of $q^{\alpha^{*}}$ is minimized among all possible rewritten queries of $q$.
The query rewrite problem has been proven to be NP-hard (Zhou et al., 2021). Traditional methods cannot select high-quality rules. To address this limitation, we advocate for utilizing LLM to select rewrite rules. To address the hallucination problem of LLMs (Huang et al., 2023), we perform an offline stage to extract query rewrite evidences, including rewrite Q&As and rewrite rule specifications, and store them as Q&A repository and rewrite rule specification repository respectively. During the online phase, given a SQL query, we retrieve relevant Q&As and rule specifications, and generate rewrite recipes, which outline how to rewrite a query using rewrite Q&As and rule specifications, assisting LLMs in comprehending the rewrite evidences. With the assistance of the rewrite recipe, LLMs are then guided through a step-by-step process to judiciously select and apply rewrite rules to the query. Next we formally define these notations.
Table 3. Rewrite Rule Specification v.s. Rewrite Rule.
Rule Specification Rewrite Rule
Mapping Single/Multiple Rules Single Rule
Source Document/Code Code in Single Engine
Clarity LLM-Readable Compiler-Readable
Selection Match Function LLM&Match Function
@@
For instance, Figure 1 shows examples of rule specification recipe and Q&A recipe. Note that the rewrite Q&A may elaborate on application of multiple rules, and thus Q&A recipes can assist in guiding rule selection by considering the interrelations among these rules. Besides, rule specification recipes, derived from various sources like documents and codes, provide complementary insights to support query rewrite, as illustrated in Table 3.
@@ 2.2. Large Language Models
Since LLMs have hallucinations (Huang et al., 2023), retrieval-augmented generation (RAG) has been proposed to mitigate this issue by indexing task-specific knowledge, retrieving relevant content for a given query, and generating answers based on the retrieved context (Gao et al., 2023; Zhang et al., 2025; Zhao et al., 2024; Li et al., 2024d). However, existing RAG techniques face limitations when applied directly to query rewrite for two primary reasons. First, there is absence of embedding methods capable of accurately evaluating the similarity between Q&As and the input query. For example, the conventional RAG technique relies on text embeddings for both the Q&A and the query, capturing only their semantic similarity while neglecting structural information pertinent to query rewrite. To address this gap, we propose a hybrid structure-semantics approach for retrieving relevant Q&As, as detailed in Section 5. Second, the multitude of rules presents a significant challenge for LLMs, as directly arranging these rules can lead to serious hallucination problems. This necessitates the development of a task-specific, step-by-step LLM algorithm that breaks down the query rewrite process into simpler, more manageable stages, as elaborated in Section 6.
@@ 3. The Overview of R-Bot
R-Bot includes an offline stage and an online stage (see Figure 2).
Offline Rewrite Evidence Preparation. This stage aims to extract rewrite Q&As from the Web and rewrite rule specifications from rewrite codes and database documents, subsequently storing them in Q&A repository and rule specification repository, respectively. Rewrite rule specifications are general and apply to multiple SQL queries, while rewrite Q&As are specific and pertain to individual queries. These resources assist in rewriting an online SQL query by guiding LLM-based selection of rewrite rules, without compromising the SQL equivalence inherently ensured by the rules.
@@
Besides, we construct efficient indexes to enhance the performance of Q&A and rule specification retrieval. We will discuss the detailed techniques in Section 4.
Online LLM-Guided Query Rewrite. Given an online SQL query, this stage retrieves pertinent Q&As and rule specifications, leverages them to rewrite the query, and offers reflections that not only complete the rewrite process but facilitate its further refinement. While LLM may have encountered data resembling rewrite evidences during pre-training, their sparse density often leads to hallucination in query rewrite. To address this, we guide LLM rewrite using the pertinent rewrite evidences prepared in a concise format.
(1) Structure-Semantics Evidence Retrieval. For an online SQL query, we retrieve relevant evidences from rewrite rule specification repository and Q&A repository.
$(i)$ Rule specification retrieval. As the rule specification has matching conditions and matching functions, we can easily retrieve the relevant rule specifications whose matching functions are satisfied by the input SQL query;
$(ii)$ Q&A retrieval. There are two types of Q&As potentially relevant to the SQL query. First, the SQL questioned in Q&A structurally matches the input SQL query. We require a structure-aware matching method to retrieve such Q&As. To this end, we propose a query structure embedding composed of $(i)$ the query template embedded by pre-trained embedding, and $(ii)$ one-hot embedding which represents the query’s matched rule specifications to assess the structural similarities. Second, Q&A semantically matches the input SQL query, such as the rewrite explanations provided in the answer section of the Q&A. However, SQL usually has no natural language rewrite explanations. To this end, we can leverage the retrieved rule specification to identify relevant Q&As based on their semantics. We propose a semantics embedding method to embed the rule specification and Q&A, and then assess their similarities based on a similarity function (e.g., $L^{2}$-distance (chr, 2024)). Lastly, to retrieve Q&As relevant to the input query both structurally and semantically, we merge the query’s structural embedding with its semantic embedding into a unified representation. We then build an embedding index for Q&A repository offline. For an online query, we utilize this index to efficiently retrieve top-$k$ relevant Q&As.
@@
$(i)$ Q&A-based rule selection and ordering. Given the rewrite rules and a sorted sequence of Q&As retrieved from the previous step, ranked by the relevance score, we select and rank the rules. We first initialize the score of each rule as 0. Then for each pair of a rule and a Q&A, we use LLM to evaluate whether they are relevant, i.e., whether the rule is applicable in light of the Q&A. If applicable, we increase the score of the rule by the score of this Q&A. Then by enumerating all the pairs of rules and Q&As, we can get the final score of each rule and rank the rules based on the final score. Since the relevance can be evaluated by LLM offline, this step can be efficiently executed by algorithms.
$(ii)$ Recipe-based rule selection and ordering. Building on the preliminary rule sequence established in the previous step, we utilize LLM to sift through and exclude any rules that do not align with the recipe. A straightforward way is to enumerate each pair of recipes returned by step (1) and rules returned by step $(2.i)$, and ask LLM to evaluate their relevance and rank the rules. However, there are two limitations. First, it may overlook the rule relevance, since a recipe may encompass multiple rules. Second, since the recipe is generated from queries and cannot be evaluated offline, assessing each pair with LLM becomes costly. To address this issue, we propose a filtering method to efficiently select the rules.
$(iii)$ Rule-based rewrite. Based on the selected rules, we input them to the query rewrite engine to rewrite the query.
@@
(3) Rewrite Reflection. It provides rewrite reflections to either further refine the query or to finalize the rewrite process. It has two reflection resources. The first involves getting the cost of the rewritten query from databases, comparing it with the cost of the query prior to its rewrite, and returning complete if the cost of the rewritten query is smaller; continue otherwise.  [^1] Database statistics play a crucial role in query optimization. However, LLM often struggles to directly understand complex database structures and intricate statistical data. To address this, we leverage the statistics indirectly through query costs, utilizing these costs to guide the LLM reflection mechanism. The second involves asking LLMs to check whether or not all the rewrite recipes are realized by the query rewrite in this step, and returning complete if yes; continue otherwise. So there are four possible reflections.
$(i)$ complete, complete: It finalizes the rewrite process and returns the rewritten query.
$(ii)$ complete, continue: It further rewrites the query by jumping to step (1) with the previously rewritten query as the input. Specifically, it starts a new round of LLM-guided query rewrite, where new rewrite evidences can be retrieved and new rules can be selected based on the previously rewritten query.
$(iii)$ continue, complete: It further refines the query by jumping to step $(2.ii)$, focusing on reordering the existing set of rules. Specifically, besides the recipes and the rules, we further input the rules actually used in the previous rewrite process, and instruct LLM to prioritize the unused rules.
$(iv)$ continue, continue: This approach integrates elements from branches $(3.ii)$ and $(3.iii)$. Initially, it defaults to branch $(3.iii)$ unless this path is revisited excessively, surpassing a predefined threshold. Under these circumstances, given that branch $(3.iii)$ has exhaustively explored the rules for the current query, the process transitions to branch $(3.ii)$. This shift initiates a new cycle, aiming to further refine the query.
Deployment at Huawei. We have deployed R-Bot at Huawei database GaussDB (Li et al., 2021; Li et al., 2024b). Initially, GaussDB identifies slow SQL queries, such as those with execution time exceeding one minute. Then, GaussDB utilizes R-Bot to rewrite the detected query. Next, if the version rewritten by R-Bot proves to be more efficient than the one rewritten by GaussDB, the system caches the query pattern using its domain-specific language (DSL). This pattern is integrated into GaussDB non-intrusively via a SQL-like plugin, which activates instantly without requiring database version updates or code changes, ensuring a seamless experience for front-end applications. During runtime, if a new query matches the cached pattern, GaussDB transparently invokes R-Bot to rewrite the query and apply the optimization. Furthermore, R-Bot has been widely used to address slow SQLs at Huawei. We have also validated R-Bot on a real-world dataset from China’s largest bank (ICBC) (see Section 7.5). The results demonstrate that R-Bot effectively optimizes real queries, significantly improving latency of 14 critical slow queries and reducing overall latency from 9.23 hours to 4.37 hours.